OpenStudy (anonymous):

??? Can someone help me to Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form

OpenStudy (radar):

I can help you with it. Do you have a specific question on the method to solve it?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the formula I have is this Y-y1=m(x-x1)

OpenStudy (radar):

You have all the information needed? Step one: Determine slope of the line you are to write the equation for. Do you know how to do that?

OpenStudy (radar):

O.K that will work, do you know how to determine m (the slope)?

OpenStudy (anonymous):

I kind of know but I have a huge confusion I dont undestand when do I have to use the formula , because there are examples problems in which they dont use the formula

OpenStudy (radar):

O.K. The slope m for your line is going to be the negative reciprocal of the slope of the line it is perpendicular to y=4x+7, the coefficient of the x term is the slope of that line, as you can see that slope is 4. What would be the negative reciprocal of 4?

OpenStudy (radar):

The reciprocal of a number is another number when multiplied equals 1. Example reciprocal of 2 is 1/2. Do you understand?

OpenStudy (anonymous):

-4 or 1/4

OpenStudy (anonymous):

-1/4

OpenStudy (radar):

Kind of combine what you have, reciprocal of 4 is 1/4 and the negative makes it -1/4

OpenStudy (anonymous):

ok

OpenStudy (radar):

O.K. now use your equation, that you said you were going to use, plug in -1/4 for m and use your given point (2, -3)

OpenStudy (radar):

Plug in the Y1 and X1 values.

OpenStudy (anonymous):

\[y-(-3)=-1/4(x-2)\]

OpenStudy (anonymous):

now I distributed y-(-3)=-1/4x-(-1/2) am I right?

OpenStudy (anonymous):

do I have to make that positive?

OpenStudy (radar):

Well let me see Your equation was (y-y1) = m(x-x1) plugging in our values (y-(-3))= -1/4(x-2) y+3=-x/4+1/2 y=-x/4 -2 1/2

OpenStudy (radar):

You could put in the same form, but you are correct. Slope intercept form is y=mx+b

OpenStudy (radar):

Sorry, but I gotta run, but you have a handle on it. good luck with them.

OpenStudy (anonymous):

Ok thank you so much for ur help